# Universal quantum control over Majorana zero modes

### v1

@@ III Chiral population transfer among Majorana zero modes
Figure 4: Four-level transition diagram with the coupling between LD and MZMs $\gamma_{1,k}$, $k=1,2,3$. Black solid lines describe the off-resonant driving fields on the transition $|e\rangle\leftrightarrow|\gamma_{1,k}\rangle$ with the detuning $\Delta_{e}(t)-\tilde{\Delta}_{k}(t)$, the Rabi-frequency $\Omega_{k}(t)$, and the phase $\varphi_{k}$. The grey dashed lines refer to the population transfer among MZMs along a clockwise or counterclockwise direction.
Beyond the large-detuning regime, we can apply the universal quantum control theory Jin and Jing (2025a) (see Appendix A for details) to realize a chiral population transfer Wang et al. (2016); Liu et al. (2020); Qi and Jing (2022) among the three MZMs $\gamma_{1,1}$, $\gamma_{1,2}$, and $\gamma_{1,3}$. As shown in Fig. 4, a LD coupled to MZM $\gamma_{1,k}$’s, $k=1,2,3$, by the local driving field with the detuning $\Delta_{e}(t)-\tilde{\Delta}_{k}(t)$, the Rabi-frequency $\Omega_{k}(t)$, and the phase $\varphi_{k}$. The full Hamiltonian can be obtained by Eq. (7), which is recalled as below
@@
$H_{I}(t)=\sum_{k=1}^{3}\Omega_{k}(t)e^{i\int_{0}^{t}[-\tilde{\Delta}_{k}(s)+\Delta_{e}(s)]ds+i\varphi_{k}}|e\rangle\langle\gamma_{1,k}|+{\rm H.c.},$ (24)
@@
According to UQC recipe for two-band systems Jin and Jing (2025b), the system dynamics can be described in the ancillary picture spanned by four ancillary basis states, which can be chosen as
@@
$|\mu_{3}(t)\rangle$ $=\cos\chi(t)|b_{2}(t)\rangle-\sin\phi(t)e^{-i\alpha_{3}(t)}|e\rangle,$
$|\mu_{4}(t)\rangle$ $=\sin\chi(t)|b_{2}(t)\rangle+\cos\phi(t)e^{-i\alpha_{3}(t)}|e\rangle.$
@@
which maintain the orthogonality with the relevant ancillary basis states, i.e., $\langle\mu_{n}(t)|b_{n}(t)\rangle=0$, $n=1,2$. The time-dependent parameters $\theta(t)$, $\phi(t)$, and $\chi(t)$ control the state populations, whereas $\alpha_{1}(t)$, $\alpha_{2}(t)$, and $\alpha_{3}(t)$ determine the relative phases among states. Equation (26) constitute a completed and orthornormal basis for the Hilbert space spanned by LD excited state and three MZMs.
Although each basis state in Eq. (26) serves as a nonadiabatic passage when it satisfies the von Neumann equation (45) with the Hamiltonian $H(t)$ (25), it is found that only $|\mu_{3}(t)\rangle$ and $|\mu_{4}(t)\rangle$ can be used to realize the chiral transfer of MZMs. Without loss of generality, we substitute $|\mu_{4}(t)\rangle$ into the von Neumann equation (45) and then obtain the conditions for the Rabi-frequencies
@@
$\times\sin\phi\cos\theta\Big{)}/\sin(\varphi_{1}+\alpha_{2}),$
@@
$\times\sin\phi\sin\theta\Big{)}/\sin(\varphi_{2}-\alpha_{0}+\alpha_{2}),$
$\Omega_{3}(t)$ $=-\left(\dot{\chi}\cos\phi-\dot{\phi}\tan\chi\sin\phi\right)/\sin(\varphi_{3}-\alpha_{1}+\alpha_{2}),$
@@
$\tilde{\Delta}_{1}(t)=-\Omega_{1}(t)\frac{\cot\chi}{\sin\phi\sin\theta}\cos(\varphi_{1}+\alpha_{2}),$ (29)
$\tilde{\Delta}_{2}(t)=-\Omega_{2}(t)\frac{\cot\chi}{\sin\phi\cos\theta}\cos(\varphi_{2}-\alpha_{0}+\alpha_{2})$
$+\dot{\alpha}_{0}\sin\chi\sin\phi\cos\theta,$
$\tilde{\Delta}_{3}(t)=-\Omega_{3}\frac{\cot\chi}{\cos\phi}\cos(\varphi_{3}-\alpha_{1}+\alpha_{2})+\dot{\alpha}\sin\chi\cos\phi,$
$\Delta_{e}(t)=\dot{\alpha}_{2}-\Omega_{1}(t)\sin\chi\sin\phi\sin\theta\cos(\varphi_{1}+\alpha_{2})$
$-\Omega_{2}(t)\sin\chi\sin\phi\cos\theta\cos(\varphi_{2}-\alpha_{0}+\alpha_{2})-\Omega_{3}\sin\chi$
$\times\cos\phi\cos(\varphi_{3}-\alpha_{1}+\alpha_{2}).$
With a appropriate choice of the parameters of $|\mu_{4}(t)\rangle$ in Eq. (26), the parameters in the lab can be reversely engineered by Eqs. (28) and (LABEL:CondDetu), then the system evolution can be precisely controlled to follow the desired passage. A complete loop of chiral population transfer of the three MZMs can be consisted of three sequential stages, each lasing identical time interval $T$. In particular, for the system initialized in the state $|\gamma_{1,1}\rangle$, the clockwise population transfer proceeds through Stage (i) $|\gamma_{1,1}\rangle\rightarrow|\gamma_{1,2}\rangle$ when $t\in[0,T]$, Stage (ii) $|\gamma_{1,2}\rangle\rightarrow|\gamma_{1,3}\rangle$ when $t\in[T,2T]$, and Stage (iii) $|\gamma_{1,3}\rangle\rightarrow|\gamma_{1,1}\rangle$ when $t\in[2T,3T]$. They are determined by the boundary conditions for the parameters $\theta(t)$, $\phi(t)$, and $\chi(t)$: $\theta(0)=\phi(0)=\chi(0)=\pi/2$, $\theta(T)=\pi$, $\phi(T)=\chi(T)=\pi/2$, $\phi(2T)=0$, $\chi(2T)=\pi/2$, and $\theta(3T)=\phi(3T)=\chi(3T)=\pi/2$. Without loss of generality, $\theta(t)$ and $\phi(t)$ for Stages (i), (ii), and (iii) of the $k$th loop, $k\geq 1$, can be set as
@@
Figure 5: Fidelity dynamics $\mathcal{F}_{n}(t)$ about various states $|n\rangle$, $n=\gamma_{1,1},\gamma_{1,2},\gamma_{1,3},e$, for two loops of (a) clockwise and (b) counterclockwise population transfers. The Rabi frequencies and the detunings of the Hamiltonian (25) are set by Eqs. (28) and (LABEL:CondDetu) with $\varphi_{1}-\alpha_{2}=\pi/2$, $\varphi_{2}-\alpha_{0}+\alpha_{2}=\pi/2$, $\varphi_{3}-\alpha_{1}+\alpha_{2}=\pi/2$ and (a) $\theta(t)$, $\phi(t)$, and $\chi(t)$ in Eqs. (30) and (31) and (b) $\theta(t)$, $\phi(t)$, and $\chi(t)$ in Eqs. (32), (33), and (31). The Rabi frequencies are in the same order as those in Fig. 2.
@@
$\theta(t)=\Phi(t)+\frac{\pi}{2},\theta(t)=\Phi(t)+\pi,$ (32)
@@
respectively. Both $\chi(t)$ and $\Phi(t)$ hold the preceding definition.
Figure 5 demonstrates the fidelity dynamics $\mathcal{F}_{n}(t)$ about the four relevant states $|n\rangle$, $n=\gamma_{1,1},\gamma_{1,2},\gamma_{1,3},e$, during the chiral population transfer. It is found that the population can propagate with a unit fidelity in both clockwise and counterclockwise manners. In Fig. 5(a) for the clockwise population transfer, the initial population on the first MZM $\gamma_{1,1}$ can be completely transferred to the second one $\gamma_{1,2}$ when $t=T$, despite the unwanted states are temporally populated during Stage (i). We have up to $\mathcal{F}_{e}\approx\mathcal{F}_{\gamma_{1,3}}\approx 0.5$ when $t=0.5T$. At the end of Stage (ii), the population transfers completely to the third MZM $\gamma_{1,3}$ at $t=2T$. And the loop is completed by the population returning to the state $|\gamma_{1,1}\rangle$ at $t=3T$. During the second loop $t\in[3T,6T]$, we have another perfect chiral transfer in the clockwise manner. Figure 5(b) demonstrates the counterclockwise population transfer with the sequences: the population in the state $|\gamma_{1,1}\rangle$ can be transferred to the state $|\gamma_{1,3}\rangle$ when $t=T$, followed by the state $|\gamma_{1,2}\rangle$ when $t=2T$, and then back to the state $|\gamma_{1,1}\rangle$ when $t=3T$. The second loop of the counterclockwise population transfer exhibits identical behavior of the first loop.
In conventional protocols for chiral transfer, both spin chirality Liu et al. (2020) and bosonic-mode chirality Qi and Jing (2022) depend on an effective time-reversal-symmetry broken Hamiltonian, from either three-body interaction or Floquet engineering. Normally the Hamiltonian is obtained by the adiabatic elimination over ancillary state or mode in the large-detuning regime, resulting in a comparatively weak effective coupling strength between interested systems. Thus, these protocols are susceptible to the environment noise due to the long running period. Moveover, they demand a precise control over the local phases of individual driving fields. In sharp contrast, our protocol enables the chiral population transfer in a nonadiabatic way. It does not invoke synthesizing an artificial magnetic field and does not require particular $\varphi_{n}$ for each driving field.
Our protocol of chiral population transfer can also be scaled to the case of $M$ MZMs described by Eq. (19). Particularly, by setting the driving frequencies as $\omega_{k}(t)=-\omega+\Delta_{e}(t)-\tilde{\Delta}_{k}(t)$ and the transition frequency $\omega\gg\Omega_{k}(t),\Delta_{e}(t),\tilde{\Delta}_{k}(t)$ for $1\leq k\leq M$, the Hamiltonian (19) can become
@@
in the rotating frame with respect to $H_{0}=-\Delta_{e}(t)|e\rangle\langle e|-\sum_{k=1}^{M}\tilde{\Delta}_{k}(t)|\gamma_{1,k}\rangle\langle\gamma_{1,k}|$, under the rotating-wave approximation. In this $1+M$-dimensional two-band model, the recipe Jin and Jing (2025b) provides two ancillary basis states, as the superpositions of the LD excited state and MZMs, either of which can be activated as a useful passage for chiral population transfer.
IV Conclusion
In the framework of universal quantum control theory, we construct robust, fast, and high-fidelity braiding operations for Majorana zero modes that are coupled to a common two-level local defect. Also, we propose a protocol to realize the perfect chiral population transfer among three MZMs. The braiding operation between an arbitrary pair of MZMs is based on the mediated exchange interaction by using the largely detuned driving fields at LD and the subsequent nonadiabatic passage. Moreover, our universal passage could employ a rapidly-varying global phase instead of extra driving fields to significantly suppress the systematic errors in LD and the driving fields on it. In the large splitting regime of the local defect, we can construct an effective Hamiltonian for the excited state of LD and the three MZMs, by which the population can propagate in both clockwise and counterclockwise manners among three MZMs with a unit fidelity. Our protocol demonstrates fault-tolerant, nonadiabatic and scalable braiding operations and the chiral population transfer with no time-reversal-symmetry broken Hamiltonian, which are fundamental and essential for the topological quantum computation and quantum information.

### v2

@@ III Chiral population transfer among Majorana zero modes
Figure 5: Four-level transition diagram with the coupling between LD and MZMs $\gamma_{1,k}$, $k=1,2,3$. Black solid lines describe the off-resonant driving fields on the transition $|e\rangle\leftrightarrow|\gamma_{1,k}\rangle$ with the detuning $\Delta_{e}(t)-\tilde{\Delta}_{k}(t)$, the Rabi-frequency $\Omega_{k}(t)$, and the phase $\varphi_{k}$. The gray dashed lines imply the population transfer among MZMs along a clockwise or counterclockwise direction.
Beyond the large-detuning regime, we can apply the universal quantum control theory Jin2025Universal (see Appendix A for details) to realize a chiral population transfer Wang2016Mesoscopic; Liu2020Synthesizing; Qi2022Chiral among the three MZMs $\gamma_{1,1}$, $\gamma_{1,2}$, and $\gamma_{1,3}$ without elimination of LD. As shown in Fig. 5, a LD coupled to MZM $\gamma_{1,k}$’s, $k=1,2,3$, by the local driving field with the detuning $\Delta_{e}(t)-\tilde{\Delta}_{k}(t)$, the Rabi-frequency $\Omega_{k}(t)$, and the phase $\varphi_{k}$. The full Hamiltonian can be obtained by Eq. (7), which is recalled as below
@@
$H_{I}(t)=\sum_{k=1}^{3}\Omega_{k}(t)e^{i\int_{0}^{t}[-\tilde{\Delta}_{k}(s)+\Delta_{e}(s)]ds+i\varphi_{k}}|e\rangle\langle\gamma_{1,k}|+{\rm H.c.}$ (29)
@@
Using the UQC recipe for two-band systems Jin2025Entangling, the system dynamics can be described in the ancillary picture by four ancillary basis states, which can be chosen as
@@
$|\mu_{3}(t)\rangle$ $=\cos\chi(t)|b_{2}(t)\rangle-\sin\chi(t)e^{-i\alpha_{3}(t)}|e\rangle,$
$|\mu_{4}(t)\rangle$ $=\sin\chi(t)|b_{2}(t)\rangle+\cos\chi(t)e^{-i\alpha_{3}(t)}|e\rangle.$
@@
which hold the orthogonality with the relevant ancillary basis states, i.e., $\langle\mu_{n}(t)|b_{n}(t)\rangle=0$, $n=1,2$. The time-dependent parameters $\theta(t)$, $\phi(t)$, and $\chi(t)$ control the state populations, whereas $\alpha_{1}(t)$, $\alpha_{2}(t)$, and $\alpha_{3}(t)$ determine the relative phases among states. Equation (31) constitutes a completed and orthornormal basis for the Hilbert space spanned by LD excited state and three MZMs.
Although each basis state in Eq. (31) serves as a nonadiabatic passage when satisfying the von Neumann equation (50) with the Hamiltonian $H(t)$ (30), it is found that only the last two can realize a perfect chiral transfer of MZMs. Without loss of generality, we substitute $|\mu_{4}(t)\rangle$ in Eq. (31) to the von Neumann equation (50) and then obtain the conditions for the Rabi-frequencies
@@
$\times\sin\phi\cos\theta\Big)/\sin(\varphi_{1}+\alpha_{3}),$
@@
$\times\sin\phi\sin\theta\Big)/\sin(\varphi_{2}-\alpha_{1}+\alpha_{3}),$
$\Omega_{3}(t)$ $=-\left(\dot{\chi}\cos\phi-\dot{\phi}\tan\chi\sin\phi\right)/\sin(\varphi_{3}-\alpha_{2}+\alpha_{3}),$
@@
$\tilde{\Delta}_{1}(t)=-\Omega_{1}(t)\frac{\cot\chi}{\sin\phi\sin\theta}\cos(\varphi_{1}+\alpha_{3}),$ (34)
$\tilde{\Delta}_{2}(t)=\dot{\alpha}_{1}-\Omega_{2}(t)\frac{\cot\chi}{\sin\phi\cos\theta}\cos(\varphi_{2}-\alpha_{1}+\alpha_{3})$
$\tilde{\Delta}_{3}(t)=\dot{\alpha}_{2}-\Omega_{3}\frac{\cot\chi}{\cos\phi}\cos(\varphi_{3}-\alpha_{2}+\alpha_{3}),$
$\Delta_{e}(t)=\dot{\alpha}_{3}-\Omega_{1}(t)\tan\chi\sin\phi\sin\theta\cos(\varphi_{1}+\alpha_{3})$
$-\Omega_{2}(t)\tan\chi\sin\phi\cos\theta\cos(\varphi_{2}-\alpha_{1}+\alpha_{3})$
$-\Omega_{3}\tan\chi\cos\phi\cos(\varphi_{3}-\alpha_{2}+\alpha_{3}).$
With Eqs. (33) and (LABEL:CondDetu), the system evolution can be precisely manipulated to follow the desired passage. A complete loop of chiral population transfer of the three MZMs can be divided into three sequential stages, each lasing an identical time interval $T$. In particular, for the system initialized in the state $|\gamma_{1,1}\rangle$, the clockwise transfer proceeds through stage (i) $|\gamma_{1,1}\rangle\rightarrow|\gamma_{1,2}\rangle$ when $t\in[0,T]$, stage (ii) $|\gamma_{1,2}\rangle\rightarrow|\gamma_{1,3}\rangle$ when $t\in[T,2T]$, and stage (iii) $|\gamma_{1,3}\rangle\rightarrow|\gamma_{1,1}\rangle$ when $t\in[2T,3T]$. They are determined by the boundary conditions for the parameters $\theta(t)$, $\phi(t)$, and $\chi(t)$: $\theta(0)=\phi(0)=\chi(0)=\pi/2$, $\theta(T)=\pi$, $\phi(T)=\chi(T)=\pi/2$, $\phi(2T)=0$, $\chi(2T)=\pi/2$, and $\theta(3T)=\phi(3T)=\chi(3T)=\pi/2$. For our purpose, $\theta(t)$ and $\phi(t)$ for stages (i), (ii), and (iii) of the $k$th loop, $k\geq 1$, can be set as
@@
Figure 6: Fidelity dynamics $\mathcal{F}_{n}(t)$ about various states $|n\rangle$, $n=\gamma_{1,1},\gamma_{1,2},\gamma_{1,3},e$, for two loops of (a) clockwise and (b) counterclockwise population transfers. The Rabi frequencies and the detunings of the Hamiltonian (30) are set by Eqs. (33) and (LABEL:CondDetu) with $\varphi_{1}+\alpha_{3}=\pi/2$, $\varphi_{2}-\alpha_{1}+\alpha_{3}=\pi/2$, $\varphi_{3}-\alpha_{2}+\alpha_{3}=\pi/2$ and (a) $\theta(t)$, $\phi(t)$, and $\chi(t)$ in Eqs. (35) and (36) and (b) $\theta(t)$, $\phi(t)$, and $\chi(t)$ in Eqs. (37), (38), and (36). The Rabi frequencies are in the same order as those in Fig. 2.
@@
$\theta(t)=\Phi(t)+\frac{\pi}{2},\quad\theta(t)=\Phi(t)+\pi,$ (37)
@@
respectively. Both $\chi(t)$ and $\Phi(t)$ hold the preceding definitions.
Figure 6 demonstrates the fidelity dynamics $\mathcal{F}_{n}(t)$ about the four relevant states $|n\rangle$, $n=\gamma_{1,1},\gamma_{1,2},\gamma_{1,3},e$, describing the chiral population transfer. It is found that the population can propagate with a unit fidelity at the ends of each stages in both clockwise and counterclockwise manners. In Fig. 6(a) for the clockwise population transfer, the initial population on the first MZM $\gamma_{1,1}$ can be completely transferred to the second one $\gamma_{1,2}$ when $t=T$, despite the unwanted states are temporally occupied during stage (i). We have up to $\mathcal{F}_{e}\approx\mathcal{F}_{\gamma_{1,3}}\approx 0.5$ when $t=0.5T$. At the end of stage (ii), the population transfers completely to the third MZM $\gamma_{1,3}$ at $t=2T$. And the loop is completed by the population returning to the state $|\gamma_{1,1}\rangle$ at $t=3T$. During the second loop $t\in[3T,6T]$, we have another perfect chiral transfer in the clockwise manner. Figure 6(b) demonstrates the counterclockwise population transfer with the sequences: the population in the state $|\gamma_{1,1}\rangle$ can be transferred to the state $|\gamma_{1,3}\rangle$ when $t=T$, followed by the state $|\gamma_{1,2}\rangle$ when $t=2T$, and then back to the state $|\gamma_{1,1}\rangle$ when $t=3T$. The second loop of the counterclockwise population transfer exhibits identical behavior of the first loop.
In conventional protocols for chiral transfer, both spin chirality Liu2020Synthesizing and bosonic-mode chirality Qi2022Chiral depend on an effective time-reversal-symmetry broken Hamiltonian, from either three-body interaction or Floquet engineering. Normally the Hamiltonian is obtained by the adiabatic elimination over ancillary state or mode in the large-detuning regime, resulting in an effective coupling strength (typically it is weak) among the interested subsystems. Thus, these protocols are susceptible to the environment noise due to the long running period. Moveover, they demand a precise control over the local phases of individual driving fields. In sharp contrast, our protocol enables the chiral population transfer in a nonadiabatic way. It does not invoke synthesizing an artificial magnetic field and does not require a particular phase $\varphi_{n}$ for each driving field.
Our protocol for chiral population transfer can also be scaled with a linear resource overhead in case of $M$ MZMs described by Eq. (19). Particularly, by setting the driving frequencies as $\omega_{k}(t)=-\omega+\Delta_{e}(t)-\tilde{\Delta}_{k}(t)$ and the transition frequency $\omega\gg\Omega_{k}(t),\Delta_{e}(t),\tilde{\Delta}_{k}(t)$ for $1\leq k\leq M$, the Hamiltonian (19) can become
@@
in the rotating frame with respect to $H_{0}=-\Delta_{e}(t)|e\rangle\langle e|-\sum_{k=1}^{M}\tilde{\Delta}_{k}(t)|\gamma_{1,k}\rangle\langle\gamma_{1,k}|$, under the rotating-wave approximation. For this $1+M$-dimensional model, the recipe in Ref. Jin2025Entangling has provided two ancillary basis states, as the superpositions of the LD excited state and MZMs, either of which can be activated as a useful passage for chiral population transfer.
IV Conclusion and comparison
In the theoretical framework of universal quantum control, we construct high-fidelity, nonadiabatic, fault-tolerant, and linear-scalable braiding operations for Majorana zero modes that are coupled to a common two-level local defect. The braiding operation between an arbitrary pair of MZMs is generated on the induced exchange interaction by using the largely detuned driving fields on LD and the subsequent rotating wave approximation. Our universal passage can employ a rapidly-varying global phase rather than extra driving fields to significantly suppress the systematic errors on LD and from quasiparticle poisoning. We also propose a protocol to realize the perfect chiral population transfer among three MZMs through their direct coupling with LD. An effective Hamiltonian can be constructed for the excited state of the local defect and the three MZMs, by which the population can propagate in both clockwise and counterclockwise manners among three MZMs with a unit fidelity. Our protocol demonstrates both braiding operations and chiral population transfer without a time-reversal-symmetry broken Hamiltonian, which are fundamental and essential elements for the topological quantum computation and quantum information processing.
It is clear to observe the discrepancies between the existing control protocols for Majorana zero modes and ours. The adiabatic braiding operations can be realized through projective measurements Knapp2016Nature and Floquet engineering Bauer2019Topologically. Although a near-unit fidelity is always achievable under ideal situations, the infidelity for both Floquet-engineering and measurement-based braiding operations inversely scales with the square of control period, i.e., $(1-\mathcal{F})\sim\mathcal{O}(T^{-2})$. The measurement-based braiding operation seems faster than the Floquet-engineering protocol since the effective coupling strength between MZMs of the former is about one order higher than that of the latter. However, it demands multiple projective measurements to perform the braiding operation, and then entails a significant overhead due to the probabilistic outcome. Our protocol takes advantage in speed since in principle all the nonadiabatic passages for constructing braiding operations can be implemented within a short period constrained only by the system characteristic frequency. While all the adiabatic control is much slower than the time-variation of the system parameters. The overhead of our protocol lies mainly in the control over the hopping and pairing couplings between the mediator LD and the neighboring fermions in the Kitaev chains and the correction operations performed to suppress the systematic errors on LD. The shortcut-to-adiabaticity version of braiding operation Karzig2015Shortcuts requires direct and time-dependent exchange interactions among MZMs, which is beyond the scope of the current technique. A recent nonadiabatic braiding operation Yu2025Nonadiabatic is based on the parallel-transport condition and is therefore fragile under the systematic errors. In comparison to both adiabatic protocols and existing nonadiabatic protocols on MZMs, our protocol only exploits the time-dependence of the local driving fields on an ancillary two-level system (local lattice defect), does not involve direct manipulation over MZMs, and is widely adaptable to other control tasks, such as chiral transfer.